OpenStudy (anonymous):

Sketch the graph and find the area under the curve defined by the parametric equations x= tan^-1t y= t^3, for 0<=t<=1

OpenStudy (amistre64):

since tan(x) = t y=(tan(x))^3

OpenStudy (amistre64):

when t=0, x=0 when t=1, x=pi/4

OpenStudy (amistre64):

tan^2 tan tan^2 = sec^2-1 sec^2(tan) - tan sec(sectan) ints up to 1/2 sec^2 -tan ints up to ln|cos| i believe

OpenStudy (amistre64):

so; we should have: 1/2 sec^2(x) + ln|cos(x)| from 0 to pi/4 as a solution if i see it right

OpenStudy (anonymous):

you lost me here sec(sectan) ints up to 1/2 sec^2

OpenStudy (amistre64):

i wonder if thats the same as integrating r r = <tan-1(t), t^3> R = < ?? ,t^4/4> from 0 to 1 I cant recall in itegral for tan-1 at the moment :)

OpenStudy (amistre64):

yeah 1/2 sec^2 derives to: 2/2 sec * sec tan

OpenStudy (amistre64):

therefore, sec sectan ints up to 1/2 sec^2

OpenStudy (anonymous):

oh ok you took the integral

OpenStudy (amistre64):

yep :)